Tohban Report 2014-11-26
From RHESSI Wiki
|Start Date:||19 Nov 2014|
|End Date:||26 Nov 2014|
|List all reports|
Solar activity was low this week with only C-class flares. There are currently 8 active regions on disk. Active region 12217 is growing in the east side of the disk and is expected to produce more C-class flares.
How many GOES flares occurred?
Flares above B, C, M, X class were 0 41 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 0 28 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 0 11 0 0
There were RHESSI flares/GOES flares 229 / 41 over the time range 18-Nov-14 25-Nov-14
The SSR was emptying by the end of the daily pass set during most of the week. The maximum fill level this week was 20%.
At the last pass, the cold tip temperatures were 106.3 and 104.3 K. The cold plate temperatures were 127.4 and 125.8 K.
The spacecraft is operating nominally. Attenuator motion nominal and crycooler is also nominal. Spin rate is good. The decimation was changed from Active/Vigorous to Normal/Vigorous at 2014-324-17:32:09 UTC.
Several short data gaps occured during a night time (14:30-15:00 UTC) of 2014-11-21. Each gap was a few minutes long.
Detector 2 continue showing light leek issues.
There were no threshold changes this week.
A spike similar to what was reported last week occurred again on 2014-11-23 at around 01:40 UTC.
David Smith noted that spikes could be due to bad packets and that he will check some more details.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/-4 minutes|
|Night time data (rears)||full nighttime events|
|Require extra passes?||No|
|Requirement for moving pointer?||No|
|Detector problems?||Detector 2 light leak.|